# Large Language Models Assume People are More Rational than We Really are

### v3

@@ Acknowledgements
Experiments were supported by Azure credits from a Microsoft AFMR grant. This work and related results were made possible with the support of the NOMIS Foundation. Special thanks to Theodore R. Sumers, Jian-Qiao Zhu and Mengzhou Xia for providing invaluable feedback for this project.
@@ Appendix A Forward Modeling prompts
A person is presented with two gambling machines, and makes a choice between the machines with the goal of maximizing the amount of dollars received.
The person will get one reward from the machine they choose. A fixed proportion of 10% of this value will be paid to the participant as a performance bonus.
@@
{} people are presented with two gambling machines, and each person makes a choice between the machines with the goal of maximizing the amount of dollars received.
Each person will get one reward from the machine they choose. A fixed proportion of 10% of this value will be paid to the participant as a performance bonus.
@@
There are two gambling machines, A and B. You need to make a choice between the machines with the goal of maximizing the amount of dollars received.
You will get one reward from the machine that you choose. A fixed proportion of 10% of this value will be paid to you as a performance bonus.
@@ Appendix B Forward Modeling Results with Different Prompts
We also observed that even when LLMs are asked to act as human participants in making decisions, their outcomes remain consistently more rational than actual human behavior under CoT prompting. Table 8 shows the results of LLMs’ decisions as individual human participants compared to the maximum expected value. Both GPT-4-Turbo and GPT-4o exhibit a high correlation with the maximum expected value, each with a correlation coefficient of 0.91 and 0.92 and the MSE of 0.045 and 0.037, while still maintaining a moderate correlation with actual human behavior, as shown in Table  7.
@@ Appendix D Comparing LLMs to a Wider Range of Behavioral Models
Appendix E Full Results for Inverse Modeling
@@
Appendix F Inverse Modeling Prompts
An example inverse modeling prompts is shown in Table 10. First, the context of the experiment is introduced, then the choices are listed, and lastly the LLM is asked to reply with which comparison more strongly suggests that the decision-maker prefers a certain target item.
@@ Appendix F Inverse Modeling Prompts
Prompts for Inverse Modeling:
@@
Please respond with either "Choice 1" or "Choice 2". Do no include anything else in your answer.
Appendix G 47 Decisions used in Inverse Decision-Making Experiment

### v4

@@ Acknowledgements
Experiments were supported by Azure credits from a Microsoft AFMR grant. This work and related results were made possible with the support of the NOMIS Foundation. Special thanks to Theodore R. Sumers, Jian-Qiao Zhu and Mengzhou Xia for providing invaluable feedback for this work.
@@ Appendix A Forward Modeling prompts
Prompts for forward modeling zero-shot are in Table 5, and prompts for forward CoT are in Table 6.
@@
A person is presented with two gambling machines, and makes a choice between the machines with the goal of maximizing the
amount of dollars received.
The person will get one reward from the machine they choose. A fixed proportion of 10% of this value will be paid to the
participant as a performance bonus.
@@
{} people are presented with two gambling machines, and each person makes a choice between the machines with the goal of
maximizing the amount of dollars received.
Each person will get one reward from the machine they choose. A fixed proportion of 10% of this value will be paid to the
participant as a performance bonus.
@@
There are two gambling machines, A and B. You need to make a choice between the machines with the goal of maximizing the
amount of dollars received.
You will get one reward from the machine that you choose. A fixed proportion of 10% of this value will be paid to you as a
performance bonus.
@@
Table 6: CoT prompting for forward modeling paradigms
LLMs predict individual human participant’s choice:
A person is presented with two gambling machines, and makes a choice between the machines with the goal of maximizing the
amount of dollars received.
The person will get one reward from the machine they choose. A fixed proportion of 10% of this value will be paid to the
participant as a performance bonus.
If the reward is negative, their bonus is set to \$0.
Machine A: {}
Machine B: {}
Which machine does the person choose?
Let’s think step by step before answering with A or B:
LLMs predict the human choice distribution:
{} people are presented with two gambling machines, and each person makes a choice between the machines with the goal of
maximizing the amount of dollars received.
Each person will get one reward from the machine they choose. A fixed proportion of 10% of this value will be paid to the
participant as a performance bonus.
If the reward is negative, their bonus is set to \$0.
Machine A: {}
Machine B: {}
How many people choose Machine A?
How many people choose Machine B?
Let’s think step by step before providing the final output.
Please provide the percentage of people who choose Machine A and Machine B in the json format.
LLMs act as human participant:
There are two gambling machines, A and B. You need to make a choice between the machines with the goal of maximizing the
amount of dollars received.
You will get one reward from the machine that you choose. A fixed proportion of 10% of this value will be paid to you as a
performance bonus.
If the reward is negative, your bonus is set to \$0.
Machine A: {}
Machine B: {}
Which machine do you choose?
Let’s think step by step before answering with A or B:
@@ Appendix B Forward Modeling Results with Different Prompts
We also observed that even when LLMs are asked to act as human participants in making decisions, their outcomes remain consistently more rational than actual human behavior under CoT prompting. Table 10 shows the results of LLMs’ decisions as individual human participants compared to the maximum expected value. Both GPT-4-Turbo and GPT-4o exhibit a high correlation with the maximum expected value, each with a correlation coefficient of 0.91 and 0.92 and the MSE of 0.045 and 0.037, while still maintaining a moderate correlation with actual human behavior, as shown in Table  9.
@@ Appendix D Comparing LLMs to a Wider Range of Behavioral Models
Appendix E Ablation Study for Forward modeling
E.1 Different temperatures
Originally, we chose temperature = 1 for some tasks and temperature = 0.7 for others. While 1 is the default temperature used by most models, 0.7 is another common option, used by works such as AlpacaEval (Li et al., 2023) on instruction following tasks.
To further validate the robustness of our findings, we test the forward modeling experiment using GPT-4o with temperatures 0.0, 0.5, 1.0, 1.5, and 2.0 for both zero-shot and CoT conditions, and compute spearman correlation with both human and max expected value, shown in Tables 12 and 13. We find that varying temperature does not substantially alter our findings other than when conducting CoT with high temperatures, where models become unable to circle back to an answer due to noisy outputs.
Table 12: Comparison of Zero-shot vs. Max Expected Value and CoT vs. Max Expected Value across temperature settings for GPT-4o.
Zero-shot vs. Max Expected Value CoT vs. Max Expected Value
Temp Spearman Temp Spearman
0.0 0.313 0.0 0.540
0.5 0.272 0.5 0.515
1.0 0.284 1.0 0.539
1.5 0.302 1.5 0.351
2.0 0.308 2.0 model unable
Table 13: Comparison of Zero-shot vs. Human and CoT vs. Human across temperature settings for GPT-4o.
Zero-shot vs. Human CoT vs. Human
Temp Spearman Temp Spearman
0.0 0.533 0.0 0.540
0.5 0.381 0.5 0.515
1.0 0.341 1.0 0.518
1.5 0.489 1.5 0.528
2.0 model unable 2.0 model unable
E.2 Demographic Groups for Decision Makers
To alleviate concerns about the simplicity of the experimental setup and further validate our findings, we conducted a follow-up experiment replicating our original analysis on various demographic profiles with varying age and gender (Tables 14 and 15). We find that our analyses stay consistent across all these factors, suggesting that our findings are at least somewhat robust to variations in context and prompt, as well as differing profiles of people.
E.3 Less Rational Decision Makers
For additional prompt variation, we also prompt the LLM to make the same predictions for a monkey rather than a human (Tables 14 and 15, bottom row). When performing both zero-shot and CoT, models accurately reflect that monkeys are less rational and the resulting predictions are less correlated with max expected value. However, for the CoT condition, the correlations are also increased with actual human behavior. This supports our conclusions with an example where the model predicts less rational decisions, but these are actually more in-line with how humans actually behave.
Table 14: Comparison of Zero-shot vs. Max Expected Value and CoT vs. Max Expected Value.
Zero-shot vs. Max Expected Value CoT vs. Max Expected Value
Demographic Spearman p-value Demographic Spearman p-value
Woman 18-35 0.289 9.44E-21 Woman 18-35 0.970 <1E-150
Woman 45-60 0.313 4.23E-24 Woman 45-60 0.972 <1E-150
Woman 65-85 0.285 9.91E-21 Woman 65-85 0.969 <1E-150
Man 18-35 0.335 2.50E-27 Man 18-35 0.971 <1E-150
Man 45-60 0.308 3.37E-23 Man 45-60 0.965 <1E-150
Man 65-85 0.348 3.98E-29 Man 65-85 0.964 <1E-150
Monkey 0.208 3.49E-13 Monkey 0.867 7.37E-123
Table 15: Comparison of Zero-shot vs. Human and CoT vs. Human.
Zero-shot vs. Human CoT vs. Human
Demographic Spearman p-value Demographic Spearman p-value
Woman 18-35 0.529 2.75E-73 Woman 18-35 0.549 9.23E-80
Woman 45-60 0.519 6.07E-70 Woman 45-60 0.546 3.54E-79
Woman 65-85 0.520 2.75E-73 Woman 65-85 0.550 3.70E-80
Man 18-35 0.499 4.18E-64 Man 18-35 0.528 2.10E-78
Man 45-60 0.530 1.43E-73 Man 45-60 0.539 1.34E-76
Man 65-85 0.508 8.54E-67 Man 65-85 0.531 3.57E-74
Monkey 0.482 2.18E-59 Monkey 0.630 8.47E-112
Appendix F Full Results for Inverse Modeling
@@
Appendix G Inverse Modeling Prompts
Example inverse modeling prompts for zero-shot and CoT are shown in Tables 16 and 17. First, the context of the experiment is introduced, then the choices are listed, and lastly the LLM is asked to reply with which comparison more strongly suggests that the decision-maker prefers a certain target item.
@@ Appendix G Inverse Modeling Prompts
Inverse Modeling, zero-shot:
@@
Please respond with either "Choice 1" or "Choice 2". Do not include anything else in your answer.
Table 17: Example prompt for inverse modeling, chain-of-thought.
Inverse Modeling, CoT:
Choice 1 was made between the following bags:
Bag 1: red, brown, yellow, blue.
Bag 2: black.
The person making the choice chose Bag 2.
Choice 2 was made between the following bags:
Bag 1: yellow, black, red, brown.
The person making the choice chose Bag 1.
Let’s think step by step.
Appendix H 47 Decisions used in Inverse Decision-Making Experiment